OpenStudy (anonymous):

h(x)=[3x-1] i need help plugging in h(2.5)

OpenStudy (saifoo.khan):

h(x) = [3x-1] h(2.5) = [3(2.5) -1] h(2.5) = [7.5 -1] h(2.5) = [6.5 ] answer

OpenStudy (anonymous):

ok i get it now